OpenStudy (anonymous):

Given the normal distribution with mean 100 and standard deviation 10, find the number of standard deviations the measurement is from the mean. Express the answer as a positive number in decimal form. Measurement = 139

OpenStudy (anonymous):

\[ z = \frac{x-\mu}{\sigma} = \frac{139-100}{10} =\ldots \]

OpenStudy (anonymous):

129?

OpenStudy (anonymous):

The distance from the mean is just: \(|139-100| = |39|=39\). You need to divide it by the standard deviation \(10\), to get it into terms of the number of standard deviations away it is.